supposen the second charge in the previous problem is -1*10^-6c.Locate the position where a third charge will not expierience a net force.

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Pallavi,

As one charge is + and other is -ve definitely the third charge should be negative as the negative charge has less magnitude compared to + ve charge.

third charge should be on the the line joining the two charges.

it should be away from + ve and closeer to - ve.

it sholud not be in between the two.

then let third charge be x distance from -ve charge then it will be at 10+x from + charge.

|F1|=|F2|

2*q/(10+x)2=1*q/(x)2

2x2=100+x2+20x

x2-20x-100=0

x=10±10√2 = 10+10√2